OpenStudy (anonymous):

@mathstudent55 @thomaster I want to know if I am doing this correctly? 2+5i/3i then i multiply that by -3i/-3i = 6i-15i/-9i^2 am i done?

OpenStudy (anonymous):

I think you mean to ask this: \[\frac{2+5i}{3i} \times \frac{-3i}{-3i}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well what is \[\frac{-3i}{-3i}=?\]

OpenStudy (mathstudent55):

What are you trying to do? What are the instructions with the problem?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@1DEA

OpenStudy (anonymous):

@1DEA

OpenStudy (mathstudent55):

If you are trying to write the complex number in standard form, then you can do it more easily. All you need to do is multiply the fraction by i/i Multiplying the fraction by -3i/(-3i) will also get you to the correct answer, but it will involve more work.

OpenStudy (mathstudent55):